OpenStudy (anonymous):

what is the constant term that completes the trinomial x^2-18x also I want to know how you got it

OpenStudy (kohai):

You can complete the square for this, I believe

OpenStudy (kohai):

To do that you have to divide the middle term (-18) by 2 and then square it. -18/2 = 9 9^2 = 81

OpenStudy (anonymous):

any number will make a "trinomial" perhaps you are attempting to complete a square?

OpenStudy (anonymous):

are you looking for a perfect square? in which case you'd want the constant to be 81 essentially when you have a quadratic in the form: a*x^2 + b*x + c and you know b (we'll just keep "a" as 1 in this case to keep it simple) c has to equal (b/2)^2 Thats because when you factor the quadratic to a form of (x+d)^2 where d is a constant, it expands into x^2 + 2d*x + d^2 now we can again compare this to our original and equate them to solve for our unknowns original = new x^2 + b*x + c = x^2 +2*d*x + d^2 the x^2 terms cancel, and you are left with b = 2d and c = d^2 You know b is -18, so now you can solve for d, and then c. d = b/2 = -18/2 = -9 c=(-9)^2 = 81 so 81 is your constant term

OpenStudy (anonymous):

Thanks!